OpenStudy (anonymous):

what is the inverse function? F(x)=x/2+7 A. F^-1(y)=7(y+2) B. F^-1(y)=2(y+7) C. F^-1(y)=2(y-7) D. F^-1(y)=7(y-2)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

c

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Let y=f(x). Where the original function sends x to y, the inverse function will send y to x. Then we rewrite the original function \[y=\frac{x}{2}+7\]and swap the variables\[x=\frac{y}{2}+7\]and solve for y\[x-7=\frac{y}{2}\]\[2(x-7)=y\]and that's your inverse function.

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

told cha lol

OpenStudy (anonymous):

D

OpenStudy (anonymous):

it's C.

OpenStudy (anonymous):

CCCCCCCCCCCCCCCCCCCCCCCCCCCCC

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

:(:(:(:(:(:(:(:(:(:(

OpenStudy (anonymous):

:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(:(

Directrix (directrix):

@MathDude000 @sourwing @cunningabe Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct